# SourceNet: Interpretable Sim-to-Real Inference on Variable-Geometry Sensor Arrays for Earthquake Source Inversion

### v1

@@ 7 Theoretical Implications
Amortized Inference vs. Classical Optimization. Classical geophysical inversion treats every event as a new optimization puzzle ($O(K)$), which makes it sensitive to initialization and computationally expensive. SourceNet represents a paradigm shift towards amortized inference (Gershman & Goodman, 2014). By paying the heavy computational cost upfront during training, the model learns the global inverse operator to allow instantaneous ($O(1)$) prediction (Adler & Öktem, 2017; Ongie et al., 2020). Moreover, unlike deterministic solvers that fit a single solution, the model learns a posterior mean that naturally marginalizes out the random uncertainties (e.g. velocity model errors) introduced during training, which leads to robustness by design.
Self-Attention as Data-Driven Green’s Functions. Mathematically, the Set Transformer implements a form of learned kernel regression (Tsai et al., 2019), so that its self-attention mechanism computes a similarity matrix that somewhat acts as a data-driven Green’s function (Kovachki et al., 2023). Instead of relying on analytical wave equations, which can be too ideal, the model learns to weigh sensor contributions based on their signal data and geometric leverage. So physically, it perform beamforming by dynamically propagating information on the observations that maximize information gain.
@@ 8 Conclusion
In this work, we proposed SourceNet, a framework that bridges deep learning and rigorous physical inference on variable-geometry arrays. By treating sensor arrays as flexible sets and bridging the Sim-to-Real gap through physics-structured randomization, we demonstrated that AI can resolve the non-linear complexities of wave propagation without relying on massive labeled real-world datasets. Our findings offer two broader insights for the AI for Science community. First, architecture should follow physics. The success of our Set Transformer over pooling baselines confirms that capturing relational dependencies (e.g., relative phases) is theoretically needed for wavefield inversion. Second, interpretability leads to discovery. SourceNet proved to be more than a predictor, as it acted as a scientific agent as it autonomously discovers the geometry gap of the sensor network and recovers optimal experimental design from data alone. By enabling models to reason over sparse, irregular observations, we address a bottleneck in scaling AI to the physical world. Towards physical Foundation Models, the ability to learn invariant operators from randomized physics, and to diagnose the limitations of the observing system itself, will be critical for turning data to scientific discovery.

### v2

@@ 7 Theoretical Implications
Amortized Inference vs. Classical Optimization. Classical geophysical inversion treats every event as a new optimization puzzle ($O(K)$), which makes it sensitive to initialization and is computationally expensive. SourceNet represents a paradigm shift towards amortized inference (Gershman & Goodman, 2014). By paying the heavy computational cost upfront during training, the model learns the global inverse operator to allow constant ($O(1)$) prediction (Adler & Öktem, 2017; Ongie et al., 2020). Moreover, unlike deterministic solvers that fit a single solution, the model learns a posterior mean that naturally marginalizes out the random uncertainties (e.g. velocity model errors) introduced during training, which leads to robustness by design, a property that cannot be learned if running optimization for each event anew (one could specify the size of errors and add a cost in the classical case, but setting the parameters of cost function is by no means a solved problem).
Self-Attention as Data-Driven Green’s Functions. Mathematically, the Set Transformer functions as a learned kernel regression (Tsai et al., 2019), so that its self-attention mechanism computes a similarity matrix that acts as a data-driven Green’s function (Kovachki et al., 2023). Instead of relying on idealized analytical wave equations, the model learns to weigh sensor contributions based on their signal data and geometric leverage. So physically, it performs beamforming by dynamically propagating information on the observations that maximize information gain.
@@ 8 Conclusion
In this work, we proposed SourceNet, a framework that bridges deep learning and rigorous physical inference on variable-geometry arrays. By treating sensor arrays as flexible sets and bridging the Sim-to-Real gap through physics-structured randomization, we demonstrated that AI can resolve the non-linear complexities of wave propagation without relying on massive labeled real-world datasets. Our findings offer two broader insights for the AI for Science community. First, architecture should follow physics. The success of our Set Transformer over pooling baselines confirms that capturing relational dependencies (e.g., relative phases) is theoretically needed for wavefield inversion. Second, interpretability leads to discovery. When inspecting its parameters, SourceNet proved to be a scientific agent that autonomously discovers the geometry of the sensor network and learns optimal experimental design without manual (human) specification. By enabling models to reason over sparse, irregular observations, we address a bottleneck in scaling AI to the physical world. In the path for physical foundation models, the ability to learn invariant operators from randomized physics, and to diagnose the limitations of the observing system itself, will be critical for discovering scientific truths in real-world data.
Impact Statement
This paper presents work with goal of advancing the field of Machine Learning for Earth Science. Our method improves the monitoring of seismic hazards, and would potentially aid in earthquake early warning and rapid response. There are no significant ethical concerns regarding surveillance or discrimination associated with this work. However, as with all deep learning models, the energy consumption for training (on 100k synthetic events) should be acknowledged, though our efficient fine-tuning protocol minimizes this footprint compared to training from scratch.
Appendix A Implementation Details and Hyperparameters
To ensure full reproducibility, we provide the specific architectural specifications and training hyperparameters used in SourceNet. The framework was implemented in PyTorch using a modular design to support dynamic configuration management (Hydra). All experiments were executed on NVIDIA A100 (40GB) GPUs.
A.1 Model Architecture
SourceNet adopts a hybrid architecture combining CNNs for local waveform feature extraction and Transformers for global event aggregation. The model consists of approximately 1.5M trainable parameters. This lightweight design enables high-throughput inference, making it suitable for real-time monitoring pipelines.
Table 2: SourceNet Hyperparameters. Values are chosen based on grid search validation performance.
Parameter Value
Station Encoder (Siamese 1D-ResNet)
Input Channels 6 (P-wave) + 6 (S-wave)
Scalar Feature Dim 20
Conv Kernel Size 7
ResNet Blocks 3
Station Embedding Dim ($d_{model}$) 128
Event Aggregator (Transformer)
Layers 3
Attention Heads 4
Feedforward Dimension 256
Dropout Rate 0.1
Positional Encoding None (Permutation Invariant)
Stage 1: Synthetic Pre-training
Dataset Size 100,000 synthetic events
Batch Size 512
Optimizer AdamW (Loshchilov & Hutter, 2017)
Learning Rate $2\times 10^{-4}$
Weight Decay $1\times 10^{-5}$
Loss Function MSE Loss
Early Stopping Patience 30 epochs
Stage 2: Real-world Fine-tuning
Dataset Size $\sim$1,700 training events (SCSN)
Batch Size 64
Optimizer AdamW
Learning Rate (Backbone) $2\times 10^{-6}$
Learning Rate (Heads) $2\times 10^{-6}$
Weight Decay $1\times 10^{-4}$
Loss Function Focal L1 Loss ($\gamma=1.5,\beta=1.0$)
Sampling Strategy WeightedRandomSampler (Balancing MT components)
Early Stopping Patience 50 epochs
A.2 Training Strategy Details
Dynamic Batching: Unlike standard fixed-size inputs, SourceNet handles a variable number of seismic stations per event. We implement a dynamic collate function that pads batches to the maximum station count within the current batch (masked during attention computation), optimizing GPU memory usage.
Two-Stage Transfer Learning:
  1. 1.
    Physics-Informed Pre-training: We first train the model on a large-scale synthetic dataset generated using Green’s functions (fk method) with randomized 1D velocity models. This forces the model to learn the fundamental physics of wave propagation and source mechanics.
  2. 2.
    Sim-to-Real Fine-tuning: We then fine-tune the model on a smaller set of high-quality real earthquake recordings from Southern California. A lower learning rate ($2\times 10^{-6}$) is used to prevent catastrophic forgetting of the physics priors, while Focal Loss is introduced to focus on hard examples (rare focal mechanisms).
Appendix B Performance on Synthetic Domain: Establishing the Theoretical Upper Bound
We evaluate the model on a held-out test set of synthetic events generated using the same PSDR engine but with unseen source parameters and geometries.
Figure 6 illustrates the training dynamics and regression performance on the synthetic domain. The learning curve (Figure 6a) shows smooth convergence without overfitting, indicating that the model successfully captures the complex mapping from waveforms to source parameters under the physics-structured randomization. The parity plots (Figure 6b) reveal near-perfect linearity for both Moment Magnitude ($M_{w}$) and Moment Tensor components ($M_{ij}$). The Mean Absolute Error (MAE) for magnitude is extremely low (0.066), confirming that the scalar tower correctly encodes geometric spreading.
Figure 7 presents the error distribution of the focal mechanism solutions. The model achieves a mean Kagan angle of $6.6^{\circ}$ and a median of $5.0^{\circ}$. This implies that in a controlled physical environment (where wave propagation is fully described by the velocity models provided), SourceNet can recover the source mechanism with near-perfect precision. The gap between this synthetic performance ($6.6^{\circ}$) and the real-world performance ($23.9^{\circ}$) quantifies the residual ”Reality Gap” caused by unmodeled factors such as complex site effects and unknown 3D velocity heterogeneities.
Figure 8 provides a visual comparison of the predicted vs. ground truth focal mechanisms (”beachballs”) for held-out events. The visual agreement is consistent across various faulting types (strike-slip, normal, reverse), confirming that the model does not suffer from mode collapse.
Figure 6: Synthetic Pre-training Performance. (a) Training and validation loss curves (MSE) over 150 epochs. (b) Parity plots for Magnitude ($M_{w}$) and Moment Tensor components ($M_{xx},M_{yy},\dots$) on the synthetic test set. The tight alignment along the $y=x$ diagonal confirms the model’s capacity to invert the physics-based simulations.
Figure 7: Error Distribution on Synthetic Data. Histogram of Kagan Angle errors. The extremely low mean error ($6.6^{\circ}$) and median ($5.0^{\circ}$) demonstrate the high fidelity of the architecture in the absence of domain shift.
Figure 8: Visualizing Focal Mechanisms (Synthetic). A grid comparison of ground truth (black) vs. SourceNet predicted (blue) beachballs for a random subset of synthetic test events. The visual agreement confirms accurate recovery of fault plane orientations.
Appendix C Real-World Fine-tuning Dynamics and Qualitative Results
Following pre-training, we fine-tune SourceNet on a small set of real-world events ($\sim$1,700 training samples) from the Southern California catalog.
Figure 9a shows the learning curve during fine-tuning. Despite the small dataset size, the validation loss decreases steadily and stabilizes, confirming that the invariant features learned during PSDR pre-training effectively prevent overfitting. Figure 9b shows the parity plots for real data. While there is more scatter compared to the synthetic domain (due to real-world noise and labeling uncertainty), the predictions remain unbiased and linear.
Figure 10 provides a qualitative comparison for held-out real-world events. SourceNet produces solutions that closely match the high-quality manual catalog (Yang et al., 2012), demonstrating robust Sim-to-Real transfer.
Figure 9: Real-World Fine-tuning Performance. (a) Learning curves during the fine-tuning stage. (b) Parity plots on the held-out real-world test set. Note that Magnitude ($M_{w}$) prediction remains robust ($MAE=0.11$) despite the significant domain gap in waveforms.
Figure 10: Visualizing Focal Mechanisms (Real World). Comparison of manual catalog solutions (Black) vs. SourceNet predictions (Blue) for real earthquake events. The model successfully recovers complex mechanisms even in the presence of real-world noise and sparse coverage.
Appendix D Detailed Ablation Visualizations
In the main text (Section 5.3), we summarized the impact of removing key architectural components using statistical metrics. Here, we provide detailed visualizations to qualitatively understand the failure modes of these ablated variants.
D.1 Impact of Scalar Tower (Multi-Modal Fusion)
We trained a variant of SourceNet without the Scalar Tower (i.e., using only waveforms as input). As shown in Figure 11, removing the scalar metadata causes a catastrophic collapse in Magnitude prediction. The $M_{w}$ scatter plot (top left) shows high variance and poor correlation compared to the full model (Figure 6). Waveform shapes alone are scale-ambiguous due to the trade-off between source energy and geometric spreading (distance). Explicit scalar metadata is physically essential for constraining the energy scale.
Figure 11: Ablation: w/o Scalar Tower. Parity plots for the model trained without scalar metadata. Note the significant degradation in Magnitude ($M_{w}$) prediction (MAE $0.07\to 0.18$), confirming the necessity of multi-modal fusion.
D.2 Impact of Self-Attention (Relational Reasoning)
We trained a variant where the Set Transformer was replaced by a DeepSets-style architecture (independent processing followed by global mean pooling). As shown in Figure 12, the Moment Tensor components ($M_{xx},M_{yy},\dots$) exhibit significantly higher scatter and bias compared to the full model. Pooling-based architectures fail to capture the pairwise relative phase information (e.g., polarity differences between stations) required to distinguish focal mechanisms. Self-Attention is theoretically necessary for resolving these geometric ambiguities.
Figure 12: Ablation: w/o Self-Attention (DeepSets). Parity plots for the model using simple pooling instead of attention. The estimation of Moment Tensor components degrades significantly (Mean Tensor MAE $0.06\to 0.17$), illustrating the failure to capture relational wave physics.